Assume that the (nominal) exchange rates b/w US and UK one year ago was $1.25/£ and currently the rate is $1.20/£ . Also, inflation rates during the year in US and UK were respectively 2% and 3%. Answer the following questions

a. What was the percentage change in the (nominal) value of the pound?

b. What the (nominal) exchange rate should be today if the RPPP (relative purchasing parity) holds?

c. What was the percentage change in the real exchange rate during the year?

d. How will the result in the above question affect the trade balance of US against UK?

please show steps
Answer #1

a.

Percentage\ change\frac{(1.25-1.2))}{1.25}\times 100=20\%

Pound appreciated by 20% over dollar in nominal terms

b.

if\ relative\ purchasing\ power\ parity\ holds\\ \frac{E[S_t^{d/f}]}{S_0^{d/f}}=\left [\frac{(1+E[p^d])}{1+E[p^f])}\right]^t\\ \frac{E[S_t^{d/f}]}{1.25}=\left [\frac{(1+2\%])}{1+3\%}\right ]^1\\ E[S_t^{d/f}]=\$1.24/ \pounds

c.

real\ rate_t-1=S_t^{d/f}\times \left [\frac{(1+\rho^f)} {(1+\rho^d)} \right ]\\ =1.25\times\frac{1.03}{1.02}=\$1.26/ \pounds

real\ rate_t-1=S_t^{d/f}\times \left [\frac{(1+\rho^f)} {(1+\rho^d)} \right ]\\ =1.2\times\frac{1.03}{1.02}=\$1.21/ \pounds

Therefore, percentage change is 4%

d. US goods become cheaper for UK because pound strengthened in real terms. US trade balance with UK shows a deficit due to weakening dollar against pound in real terms.